OpenStudy (anonymous):

In a right of a triangle, find the length of the side not given. b=1, c=sqrt(5)

OpenStudy (eyust707):

use a^2 + b^2 = c^2

OpenStudy (anonymous):

Assuming that sqrt5 is the hypotenuse, the other side is 2 so that: sqrt( 1^2 + 2^2) = sqrt5 i think

OpenStudy (eyust707):

(1)^2 +b^2 = (sqrt5)^2 1 + b^2 = 5 b^2 = 4 sqrt(b^2) = sqrt(4) b = 2

OpenStudy (akshay_budhkar):

which of the sides are these? the one included in right angle?? or is any1 of them a hypotenuse?

OpenStudy (eyust707):

it is standard convention to use a and b as the legs and c as the hypotenuse